OpenStudy (anonymous):

What is the equation of the line that passes through (0, -2) and has a slope of 0? a. y = -2 b. y = 2 c. x = 2

OpenStudy (midhun.madhu1987):

The equation of the line is of the form y - y1 = m(x - x1) here m = 0 So, y-y1 = 0 substitute the value of y1

OpenStudy (anonymous):

These are my options, can you please help. I don't get it at all. a. y = -2 b. y = 2 c. x = 2

OpenStudy (anonymous):

its A y=-2 sorry you had to wait

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

@xorubyrosas

OpenStudy (anonymous):

thank you @mathmagic101